# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

### v2

@@ 3 Methods
We surveyed companion chatbot users recruited from CloudResearch, administering several psychological scales and measures of chatbot usage frequency and session length. This study was approved by the institution’s IRB. Using the Qualtrics sample size estimator, we roughly estimated a minimum size of 385 with a 95% confidence level and a 5% margin of error. Our criteria for participants were that they must be:
  • 18 years old or older
  • fluent in English
  • regular users of companion chatbots (using them at least once a week for at least a month).
Participants were first recruited with a screening survey, only accessible on CloudResearch if they were fluent in English and had prior experience using AI. A researcher manually read through respondents’ brief free responses regarding their experience with companion chatbots to remove participants who used chatbots only for practical purposes (i.e. assisting with coding or interview preparation) or those who did not describe any personal experiences (i.e. only describing the service). Participants were screened and recruited through multiple rounds until we reached our target pool of qualified individuals. All qualifying participants were then invited to complete the main survey. We intentionally over-recruited to account for incomplete responses. After filtering out incomplete responses from the main survey, the final sample consisted of 404 valid responses (n = 404), slightly exceeding our initial target of 400.
We conducted the following analyses:
  • Participant Demographics: We calculated descriptive statistics to create a general overview of our participants.
  • Correlation Analysis: We calculated a Spearman correlation matrix with Benjamini-Hochberg correction as a starting point for exploring relationships in our numerical variables.
  • Multiple Regression Analysis: We performed an exploratory multiple regression analysis to construct a more nuanced model that can capture relationships between variables not reflected in the correlation analysis.
  • Participant Cluster Analysis: We performed K-means clustering of our numerical data to find distinct groups of users in our data.
  • Chatbot Usage Characteristics: We calculated descriptive statistics on multiple-choice questions regarding companion chatbot usage and examined participants’ free responses, both overall and grouped by the clusters we found. Our intention was to characterize our participants overall and provide more specific characteristics and examples of responses for each cluster. We also conducted thematic analysis of participants’ reasons for initial interest, motivations for continuing usage, and topics of conversation using a stratified random sample of 15 participants per cluster (total N = 105).
These analytical approaches were designed to address our primary research questions:
  1. 1.
    To explore motivations and uses of companion chatbots (RQ1), we analyzed the descriptive statistics of usage characteristics and examined participants’ free responses.
  2. 2.
    To investigate the relationship between chatbot usage and loneliness (RQ2), we employed correlation analysis and multiple regression analysis. These methods allowed us to identify potential mediating or moderating factors in this relationship.
  3. 3.
    To identify and characterize distinct user profiles (RQ3), we conducted K-means cluster analysis on our numerical data, including usage patterns, loneliness levels, and other psychological and social factors. We then examined chatbot usage characteristics within each cluster to provide more nuanced profiles.
Figure 3: A flowchart detailing the multi-method approach employed in this study. The diagram outlines the sequence of analytical steps, including participant recruitment, survey administration, data processing, correlation analysis, multiple regression modeling, cluster analysis, and the interpretation of both quantitative results and free-text responses.
We summarize these methods in Figure 3. By combining these quantitative and qualitative approaches, we aimed to provide a comprehensive understanding of companion chatbot usage patterns, their relationship with psychological factors such as loneliness, and the characteristics of different user groups.
3.1 Survey Design
Our survey instrument was designed to investigate the relationship between chatbot usage and loneliness, as well as potential mediating factors. The survey comprised both established scales (some adapted) and custom questions developed for this study. The full survey can be found as a Word document in Supplementary Material.
3.1.1 Established Scales and Adaptations
Our scale selection was guided by two principles: (1) measuring our primary variable of interest, loneliness, with a well-validated instrument, and (2) exploring potential mediating and moderating factors through an exploratory approach. While we acknowledge that including many scales risks multicollinearity, we addressed this through statistical controls in our analysis, detailed in the Analysis section.
Our primary dependent variable measure was the UCLA Loneliness Scale (ULS-8) [Hays and DiMatteo(1987)], demonstrating high internal consistency ($\alpha$ = .84) and construct validity through correlations with depression and inverse correlations with satisfaction with life, and social support [Wiseman et al.(1995), Wu and Yao(2008)]. It was selected for its brevity while maintaining psychometric properties of the original scale.
For our exploratory investigation of potential mediating and moderating factors, we included:
  • Lubben Social Network Scale (LSNS-6) [Lubben et al.(2006)]: Measures social isolation with strong internal consistency ($\alpha$ = .83). Selected to provide an objective measure of social connection to complement the subjective measure of loneliness.
  • Multidimensional Scale of Perceived Social Support (MSPSS) [Zimet et al.(1990)]: Assesses perceived social support. Selected for strong psychometric properties ($\alpha$ = .88) and wide usage [Wang et al.(2021)], as well as the known correlation between loneliness and social support [Ozbay et al.(2007), Wang et al.(2018), Schutter et al.(2021)].
  • Brief Rosenberg Self-Esteem Scale (B-RSES) [Monteiro et al.(2022)]: Evaluates self-esteem, a known risk factor leading to and increased by loneliness [Vanhalst et al.(2013)]. The brief scale was chosen for its shorter length but good internal consistency ($\alpha$ = .89 in a general population) and strong validity comparable to the original scale [Monteiro et al.(2022)].
  • Big Five Inventory (BFI-10) [Rammstedt and John(2007)]: Evaluates personality traits based on five factors: extraversion, neuroticism (tendency to towards anxiety and other negative feelings), agreeableness (tendency to be compassionate and cooperative), conscientiousness (tendency to be self-disciplined and goal-directed), and openness to experience (tendency to be open-minded, imaginative, and insightful) [McCrae and Costa(1999)]. It correlates highly with the full BFI (mean correlation .83) [Rammstedt and John(2007)]. Selected based on previous research linking personality traits to loneliness [Buecker et al.(2020)].
  • Attitudes Towards AI Scale [Schepman and Rodway(2020)]: Measures attitudes towards artificial intelligence with strong internal consistency for both the positive ($\alpha$ = .88) and negative ($\alpha$ = .83) subscales. Used the 8-item adaptation of the 16-item positive subscale, found in [Pataranutaporn et al.(2023)].
  • Human-Chatbot Interaction Effect Scale [Xia et al.(2024)]: Measures the effect of human-chatbot interaction on human-human interaction. We included measures of the respondent’s expectations for chatbot services ($\alpha$ = .85), trust in interactions with social chatbots ($\alpha$ = .86), perceived value of substituting people with chatbots ($\alpha$ = .91), trust in interactions with people ($\alpha$ = .93), and motivation and willingness to interact with people ($\alpha$ = .94). Changed phrasing to refer to only chatbots and people rather than service providers, shortened subscales based on factor loadings, and made two versions of the interpersonal communication intention subscales, one for familiar people and one for unfamiliar people.
  • Generalized Problematic Internet Use Scale (GPIUS2) [Caplan(2010)]: A validated measure of problematic internet use with strong psychometric properties ($\alpha$ = .91 for full scale, CFI = .95) [Caplan(2010)]. We adapted six items from the scale to specifically reference chatbot use, maintaining representation of subscales measuring compulsive use, cognitive preoccupation, and negative outcomes. Existing work shows that loneliness is both a cause and effect of problematic internet use [Kim et al.(2009)], and the GPIUS2 was developed with online social interactions in mind, making it well-suited for examining potential compulsive or problematic patterns in human-AI interaction.
  • State Empathy Scale [Shen(2010)]: A validated measure capturing affective, cognitive, and associative dimensions of empathy. The scale demonstrates strong construct validity and reliability across multiple studies, with confirmatory factor analyses supporting its unidimensional structure [Shen(2010)]. Adjusted wording to be more general for application to both chatbots and people. Selected as a way to assess how participants interpersonally connect with chatbots, as research suggests people process parasocial relationships similarly to interpersonal ones [Schiappa et al.(2005)] and there may be parallels between parasocial relationships and human-chatbot relationships.
  • Interpersonal Attraction Scale [McCroskey and McCain(1974)]: Measures three dimensions of attraction: social (desire for friendship), physical (appearance), and task (competence-based). The scale’s application to mediated relationships is supported by extensive parasocial interaction research [Schiappa et al.(2005)]. We shortened each subscale to four items based on factor loadings and created two versions: one for chatbots and one for close persons (excluding physical attraction for the latter due to potential family relationships).
  • Perceived Homophily Scale [Mccroskey et al.(1975)]: Measures perceived similarity in communication, which research shows is central to both interpersonal and parasocial relationship development [Schiappa et al.(2005)]. Modified to use unidirectional Likert statements (e.g., "like me" instead of "unlike me – like me"). Selected 6 items based on factor loadings and created parallel versions for chatbots and a close person.
  • Attributional Confidence Scale [Gudykunst and Nishida(1986)]: Assesses ability to predict others’ attitudes, feelings, and behavior [Clatterbuck(1979)]. Its relevance to human-AI interaction is supported by uncertainty reduction theory’s application to parasocial relationships; that is, relationships develop as people increase their ability to predict others’ behavior [Perse and Rubin(1989)]. Shortened to 5 items, removing items irrelevant to chatbots.
  • System Usability Scale (SUS) [Brooke(1996)]: A widely-validated measure of perceived system usability demonstrating high reliability ($\alpha$ = .91) across diverse interfaces [Bangor et al.(2008)]. Adjusted phrasing to refer specifically to chatbots. Chosen for consideration of potential links between usability and perceived closeness.
We acknowledge that modifying established scales may affect their validated psychometric properties. To manage survey length while maintaining measurement integrity, we shortened scales based on factor loadings from previous studies when available. While this approach allowed us to explore multiple potential factors, we recognize this as a limitation and address it through statistical controls for multicollinearity in our analysis, Benjamini-Hochberg correction for multiple comparisons, and conservative interpretation of results from modified scales. This exploratory approach allows us to identify potential factors warranting further investigation while maintaining statistical rigor through appropriate controls.
3.1.2 Custom Survey Items
In addition to established scales, we developed custom items to capture specific aspects of chatbot usage. Each item except usage characteristics required participants to elaborate on their choices and share thoughts not captured in the multiple-choice options as a free response, as our provided options were not expected to be exhaustive.
  • Usage Characteristics: Multiple-choice questions for frequency of chatbot use (5 options ranging from daily to once a month or less plus options indicating variation) and average session length (6 options: <15 min, 15-30 min, 10 min-1h, 1h-2h, 2h-4h, 4h+). Broad options were chosen to reduce the risk of inaccuracies from recall bias.
  • Motivations and Continued Use: Multiple-choice questions with required free-text elaboration on why participants began using chatbots and why they continue to use them; the options can be seen in Figure 5 (left graphs). These options were chosen through discussion with multiple researchers and based on uses and gratifications theory [Katz et al.(1973), Tanta et al.(2014)].
  • Conversation Topics: Questions about the topics participants discuss with chatbots, with the options chosen from discussions with multiple researchers (see top right graph of Figure 5).
  • Human vs. Chatbot Preference: Several items for participants’ preferences for human or chatbot interaction in various contexts, with the above conversation topics presented as options in a matrix question.
  • Perceived Impact on Relationships: Six Likert scales assessing participants’ perceptions of how chatbot use has affected their relationships with people by leading to (1) less interaction, (2) changes in conversation topics, and (3) greater ease, (4) greater enjoyment, (5) less engagement, and (6) less enjoyment in talking to people. These were selected through discussion with multiple researchers to roughly gauge participants’ perceived relationship changes.
3.2 Correlation and Multiple Regression
Data processing and statistical analyses were performed using Python, with the statsmodels library for regression analyses and scikit-learn for data preprocessing. We used a significance level of $\alpha=0.05$ for all statistical tests, unless otherwise specified. Usage frequency and session length were converted to linear numerical scores, with multiple responses averaged. All numerical variables were standardized using StandardScaler to improve interpretability and reduce potential multicollinearity [Aiken and West(1991)].
We computed a Spearman correlation matrix [Cohen et al.(2003)] with Benjamini-Hochberg correction for multiple comparisons[Benjamini and Hochberg(1995)]. For the regression analysis, we focused on session length as the independent variable, as it showed a significant correlation with loneliness where frequency did not.
Potential mediators were identified using the corrected p-values (p < 0.05). Mediation analysis followed Baron and Kenny’s [Baron and Kenny(1986)] approach, with Sobel tests [Sobel(1982)] assessing mediation significance. We tested moderation effects by creating interaction terms and examining their significance in regression models [Aiken and West(1991)]. To address potential multicollinearity, we calculated Variance Inflation Factors (VIF), considering variables with VIF > 5 for elimination [Hair et al.(2010), O’brien(2007)]. After removing the state empathy subscales, which were redundant with the full scale, all predictors were below the threshold.
To build our regression model, we started with all identified mediators, moderators, and confounders [Greenland et al.(1999), VanderWeele and Shpitser(2013)], including socialization and neuroticism on theoretical grounds. While retaining age and sex as control variables, we used backward elimination to remove predictors with high p-values (generally p > 0.1) to make a more parsimonious and interpretable model [Cohen et al.(2003)]. We then explored adding interaction terms one at a time, retaining significant interactions. Categorical variables (area type, education, relationship status, political party, race, employment status) were tested using ANOVA but were not retained due to non-significance.
We assessed model fit using R-squared, F-statistic, and condition number. Model assumptions were checked using Q-Q plots, residuals vs. fitted plots, and influence plots [Cook and Weisberg(1982), Fox(2015)]. The summary of the final model from our exploration can be seen in Figure 13.
To examine the potential mediators of our final model, we conducted a parallel multiple mediation analysis. This analysis was performed using a bootstrapping approach with 5,000 resamples, as recommended by [Preacher and Hayes(2008)]. The analysis controlled for several covariates, including age, sex, and personality traits (BFI). Confidence intervals (95%) for the indirect effects were calculated using the percentile method.
3.3 Cluster Analysis
We employed K-means clustering analysis to identify distinct user profiles of chatbot users based on their psychosocial characteristics. These profiles were identified based on a set of features with significant correlations (p > 0.05, corr > 0.3) to key variables, as well as features relevant to chatbot usage including usage frequency and measures of empathy and attraction with a chatbot; the features were reduced to 10 principal components. After some exploration, we used Chatbot Session Length, Frequency, Loneliness (ULS-8), and Problematic Use of Chatbot (GPIUS2) for our key variables. The optimal number of clusters was determined using a combination of quantitative metrics and qualitative interpretation:
  1. 1.
    We calculated the Within-Cluster Sum of Squares (WCSS) for a range of cluster numbers (2-14) and plotted the results to visualize the elbow method [Kodinariya and Makwana(2013)].
  2. 2.
    We computed silhouette scores for the same range of clusters to assess cluster cohesion and separation [Kaufman and Rousseeuw(2009)].
  3. 3.
    We examined the interpretability and meaningfulness of the resulting clusters for different k values, considering our research questions and domain knowledge.
The silhouette scores increased slightly as the number of clusters increased from 3 to 7, dropping to a local minimum at 8 clusters. The silhouette score fluctuated without consistent increases and there was no clear bend in the elbow method, we ultimately chose a 7-cluster solution. This decision was also based on the interpretability and distinctiveness of the resulting user profiles, aligning with our objective to understand the characteristics of different groups of users.
The final clustering was performed using the K-means algorithm with k=7, implemented in Python using the scikit-learn library [Pedregosa et al.(2012)]. The resulting clusters were then characterized based on their centroids and distribution of key variables. To validate our clustering solution, we employed a multi-faceted approach:
  1. 1.
    We conducted Kruskal-Wallis tests to examine differences among clusters for key variables (Session Length, Loneliness, and Problematic Use), as the data violated normality assumptions [Kruskal and Wallis(1952)].
  2. 2.
    After confirming significant differences with Kruskal-Wallis tests, we performed Dunn’s post-hoc test with Bonferroni correction for multiple comparisons [Dunn(1964)] to identify specific significant differences between clusters on our key variables (session length, loneliness, problematic use, and frequency).
  3. 3.
    We calculated the average silhouette score to assess the quality of cluster separation [Rousseeuw(1987)].
  4. 4.
    We performed a Multivariate Analysis of Variance (MANOVA) to test for overall differences among clusters across multiple dependent variables simultaneously [Huberty and Olejnik(2007)].
  5. 5.
    We used Box’s M-test to check the assumption of homogeneity of covariance matrices for the MANOVA [Box(1949)].
These validation techniques provide a comprehensive assessment of the cluster solution’s statistical robustness and meaningful differentiation among user profiles.
3.4 Interpretation of Quantitative Results and Free-Text Responses
To provide a comprehensive understanding of our findings, we employed a two-pronged approach to interpreting the results of our quantitative analyses. First, we calculated descriptive statistics on companion chatbot usage characteristics for both our entire dataset and the clusters we identified through K-means analysis. This provided a quantitative overview of usage patterns and user profiles.
Second, we examined the free-text responses associated with multiple-choice items to complement our quantitative findings. We conducted a thematic analysis [Braun and Clarke(2006)] of three free-response items (participants’ reasons for initial interest, motivations for continuing usage, and topics of conversation) using a stratified random sample; 15 participants were randomly selected from each cluster while maintaining equal representation across clusters (total N = 105). A researcher read through the sample of free responses, developed a codebook, and extracted overall themes for each examined item. The researcher developed initial codes by identifying recurring patterns, then grouped these codes into broader themes that captured key aspects of users’ reasons for engaging with companion chatbots.
By integrating these approaches, we were able to capture both broad trends identified through our quantitative analyses and individual nuances in user experiences expressed in their own words. This method provided a richer interpretation of the data, offering insights into the motivations, perceptions, and experiences of companion chatbot users that might not be fully captured by quantitative measures alone.
@@ 4.1 RQ1: What motivates the use of companion chatbots, and what do people use them for?
Figure 4: Percentage of total responses for multiple-choice questions regarding chatbot usage patterns among participants. Top left: chatbot session length. Top right: chatbot usage frequency. Bottom left: chatbot usage timeframe. Bottom right: most used chatbot platforms.
Figure 5: Percentage of participant responses for multiple-choice questions regarding their use of companion chatbots. Top left: reasons for using companion chatbots. Top right: conversation topics with chatbots. Bottom left: motivations for continued use. Bottom right: preferred devices for companion chatbot use.
Figure 6: Codes and themes for initial and continued usage of companion chatbots from thematic analysis of a stratified random sample of participants (15 per cluster, total N = 105).
@@ 4.1.1 Motivations for initial interest and continued use of companion chatbots
Analysis of both quantitative and qualitative data revealed diverse motivations for chatbot engagement, with several key themes emerging across both initial adoption and continued use. Descriptions of these themes can be found in Figure 6.
Technology exploration emerged as the predominant motivation, manifesting quantitatively in curiosity about AI and chatbots (30.89% initial use, 20.91% continued use). Users expressed interest in testing the technology’s capabilities and limitations: “I wanted to see if it could actually emulate a real person.”
Recreation represented another major theme, encompassing both entertainment (26.35% initial, 23.69% continued) and creative activities (17.81% initial use). Users enjoyed role-playing with characters and leveraging chatbots for creative projects: “I use chatbots to practice conversations for my D&D game, creating NPC personalities and dialogue.”
The practical utility theme emerged particularly in continued use, with users valuing chatbots for task assistance and information seeking. As one user explained: “I usually use chatbots to help with ideas like dinner ideas or information I may want to find out without using a typical search engine.” This theme aligns closest with the quantitative findings showing passing time and reducing stress (12.88% continued use) as a notable factor, though there were no directly-related multiple-choice options originally.
Another theme not explicitly reflected in the initial multiple-choice options was the safe space theme, which emerged to encompass the value of chatbots as non-judgmental environments for self-expression an emotional support. One participant noted, “I sometime feel lonely and just want to be left alone, during this time I like chatting with my AI companion because I feel safe and won’t not [sic] be judged for the inadequate decisions I have made.”
Social motivations, while less prevalent, are significant for some users. These motivations encompassed themes companionship, self-improvement, and lifestyle integration. Initial use showed loneliness (12.24%), social skill development (7.24%), and relationship seeking (3.53%) as drivers, while continued use revealed ongoing needs for companionship (9.86%), emotional experiences (6.52%) and self-discovery (6.12%). Some users explicitly sought social improvement: ‘I’m autistic, so I’m not the best at social interaction, and I found that practicing with chatbots helped.” The lifestyle integration theme emerged specifically in continued use, with some users incorporating chatbots into their daily routines: “It has become part of my life and I want to continue to learn and evolve with new technology.” This suggests that for some users, chatbots transition from novelty to utility and integration over time.
Our thematic analysis enriches our understanding beyond pure statistics, revealing how different motivations interweave and evolve throughout users’ engagement with companion chatbots. While initial interest often stems from curiosity and entertainment, continued use appears more deeply rooted in practical and emotional benefits.
Figure 7: Codes and themes for topics of conversation with companion chatbots from thematic analysis of a stratified random sample of participants (15 per cluster, total N = 105).

### v3

@@ 3 Methods
We surveyed companion chatbot users recruited from CloudResearch Connect [47], administering psychological scales and measuring chatbot usage patterns. The study was IRB-approved with a target sample of 385 participants (95% confidence level, 5% margin of error). Participants were 18+ years old, English-fluent, and regular chatbot users (weekly use for at least one month). After filtering incomplete responses, our final sample consisted of 404 participants.
3.1 Analytical Approach
Our mixed-methods approach included (1) Spearman correlation analysis of numerical variables with Benjamini-Hochberg correction, (2) multiple regression exploring relationships between chatbot usage, loneliness, and potential mediating/moderating factors, (3) K-means clustering to identify distinct user profiles based on psychosocial characteristics and usage patterns, and (4) thematic analysis from a stratified random sample (N = 105) combined with generative AI-based topic extraction of the full dataset.
3.2 Survey Instrument
Our survey combined established scales and custom items:
3.2.1 Established Scales
Our primary dependent variable was measured with the UCLA Loneliness Scale (ULS-8) [48], selected for its brevity while maintaining strong psychometric properties (Cronbach’s $\alpha$ = .84). For potential mediating and moderating factors, we included the following items: Lubben Social Network Scale (LSNS-6) [49], measuring social isolation ($\alpha$ = .83); Multidimensional Scale of Perceived Social Support (MSPSS) [50], assessing perceived social support ($\alpha$ = .88); Brief Rosenberg Self-Esteem Scale (B-RSES) [51], evaluating self-esteem ($\alpha$ = .89); Big Five Inventory (BFI-10) [52], assessing personality traits (mean correlation with full BFI: .83); Attitudes Towards AI Scale [53], measuring AI attitudes (positive subscale: $\alpha$ = .88; negative: $\alpha$ = .83); Human-Chatbot Interaction Effect Scale [9], measuring effects of human-chatbot interaction on human-human interaction; Adapted Generalized Problematic Internet Use Scale (GPIUS2) [54], measuring problematic chatbot use ($\alpha$ = .91 for original scale); State Empathy Scale [55], capturing empathy dimensions with strong construct validity; Interpersonal Attraction Scale [56], measuring attraction dimensions, with versions for chatbots and close persons; Perceived Homophily Scale [57], measuring perceived similarity in communication; Attributional Confidence Scale [58], assessing ability to predict others’ attitudes and behaviors; and System Usability Scale (SUS) [59], measuring perceived chatbot usability ($\alpha$ = .91).
Where necessary, we shortened scales based on factor loadings from previous studies to manage survey length while maintaining measurement integrity. We addressed this limitation through statistical controls for multicollinearity and conservative interpretation.
3.2.2 Custom Items
We developed custom items to collect data on chatbot usage frequency and session length, as well as usage motivations, conversation topics, and perceived impacts on relationships. Most items combined multiple-choice selections with free-response elaboration, with participants providing elaboration for reasons for initially using and continuing to use chatbots, conversation topics, perceived effects of chatbot use on human relationships (combined with Likert-scale items reporting changes in human interactions), situational preferences for chatbot versus human interaction (combined with a preference matrix across different contexts), and reasons underlying these preferences.
3.3 Statistical Analyses
All analyses were performed using Python. We calculated descriptive statistics on chatbot usage characteristics for both the entire dataset and identified clusters. Usage frequency and session length were converted to numerical scores, and all variables were standardized. For correlation analysis, we computed a Spearman correlation matrix with Benjamini-Hochberg correction (Figure˜1).
Figure 1: Correlation matrix (Spearman, Benjamini-Hochberg corrected) showing significant ($p<0.05$) correlations of psychological, social, and usage characteristics. Variables are sorted into general categories, and names have been altered for readability. The original names can be seen in Figure˜6 of the Appendix.
Regression analysis focused on session length as the independent variable (as frequency showed no significant correlation with loneliness). Potential mediators were identified using corrected p-values ($p<0.05$) following Baron and Kenny’s approach [60] with Sobel tests [61]. Moderation effects were tested by examining interaction terms [62]. Variance Inflation Factors were calculated to address multicollinearity [63, 64]. Our regression model initially included all identified mediators, moderators, and confounders [65, 66]. We used backward elimination to create a more parsimonious model [67], retaining age and sex as control variables. Model fit was assessed using R-squared, F-statistic, and condition number, with assumptions checked using Q-Q plots, residuals vs. fitted plots, and influence plots [68, 69]. We conducted parallel multiple mediation analysis using bootstrapping with 5,000 resamples [70].
We then employed K-means clustering to identify distinct user profiles. Features were selected based on significant correlations with key variables ($p>0.05$, corr >0.3>0.3) and relevance to chatbot usage, then reduced to 10 principal components. The optimal number of clusters was determined using the elbow method [71], silhouette scores [72], and interpretability considerations. We chose a 7-cluster solution based on these metrics and the distinctiveness of resulting profiles. To validate our clustering, we conducted Kruskal-Wallis tests [73] followed by Dunn’s post-hoc tests with Bonferroni correction [74], calculated average silhouette scores [75], and performed MANOVA [76] with Box’s M-test [77].
3.4 Mixed-Methods Thematic Analysis
To complement quantitative findings, we employed a mixed-methods approach combining manual thematic analysis [78] with automated topic extraction, addressing limitations of purely manual analysis (limited sample coverage) and purely automated analysis (reduced contextual sensitivity) [79]. Thematic analysis was conducted on a stratified random sample (N=105, 15 per cluster), with one researcher developing codes and grouping them into broader themes. To validate themes across the complete dataset (N=404), we implemented an automated pipeline using GPT-4 for topic extraction (see Appendix for details), processing responses in batches of 300 with instructions to identify 5-10 key themes per batch, including descriptions, prevalence estimates, and representative quotes. A researcher reviewed the LLM output to assess accuracy and distinctiveness before comparing to manual themes based on descriptions and quotes; LLM themes were matched to similar manual themes or identified as subsets where appropriate, with substantial overlap in identified patterns. This triangulation strengthened confidence in our thematic findings while demonstrating how manual depth can complement automated breadth in qualitative analysis.
@@ 4.1 RQ1: What motivates the use of companion chatbots, and what do people use them for?
Figure 2: Percentage of total responses for multiple-choice questions regarding chatbot usage patterns among participants.
@@ 4.1.1 Motivations for initial interest and continued use of companion chatbots
Participants were asked separate questions about their initial interest in starting chatbot use and their reasons for continuing use, each combining multiple-choice options with free-response elaboration. Analysis of quantitative and qualitative data revealed diverse motivations, validated through automated theme extraction across the full dataset. Descriptions of these themes can be found in Figure˜10 in the Appendix.
Technology exploration emerged as the predominant motivation (30.89% initial use, 20.91% continued use), with users expressing curiosity about AI capabilities: “I wanted to see if it could actually emulate a real person.”
Recreation represented another major theme, encompassing entertainment (26.35% initial, 23.69% continued) and creative activities (17.81% initial use). Users enjoyed role-playing and creative projects: “I use chatbots to practice conversations for my D&D game, creating NPC personalities and dialogue.”
Practical utility emerged particularly in continued use, with users valuing task assistance and information seeking: “I usually use chatbots to help with ideas like dinner ideas or information I may want to find out without using a typical search engine.”
The safe space theme encompassed chatbots as non-judgmental environments for self-expression: “I sometime feel lonely and just want to be left alone, during this time I like chatting with my AI companion because I feel safe and won’t not [sic] be judged.”
Notably, while companion chatbots are often marketed as solutions for loneliness, only 12.24% of participants initially sought them for companionship (9.86% continued). Other social motivations were also less prevalent but significant for some users, including self-improvement for social skills (7.24% initial) and lifestyle integration where chatbots become part of daily routines. While initial interest often stems from curiosity and entertainment, continued use appears more rooted in practical and emotional benefits.